OpenStudy (priyar):

The coefficient of x^7 in the expansion of (1-x-x^2 +x^3)^6 is?

OpenStudy (priyar):

pls explain using multinomial theorem..

OpenStudy (priyar):

@imqwerty

imqwerty (imqwerty):

multinomial theorem says that in \(\left(a_1+a_2+a_3+....+a_m\right)^n\) the coefficient of \(a_1^{p_1}a_2^{p_2}...a_m^{p_m}\) =\(\large \frac{n!}{p_1!.p_2!...p_m!}\) where \(p_1+p_2+...p_m=n\) 1st we need to know the ways to get \(x^7\) and then find the coefficient of each case and then add them all

OpenStudy (priyar):

ok..let me try..

imqwerty (imqwerty):

okay

OpenStudy (priyar):

there 7 such cases right?

OpenStudy (priyar):

@imqwerty

imqwerty (imqwerty):

um i didn't do the calculation tho

OpenStudy (priyar):

and sorry it took so long..

imqwerty (imqwerty):

its okay ima do the calculations

OpenStudy (priyar):

there were so many possibilities...(trial and error)..

imqwerty (imqwerty):

yes its better to convert such cases into a binomial theorem applicable form

OpenStudy (priyar):

how can we convert this?

imqwerty (imqwerty):

we can write it like this- \(1-x-x^2+x^3=(x-1)^3+2(x-1)^2\)

OpenStudy (priyar):

shouldn't it be 2(x+1)^2 over there?

OpenStudy (priyar):

oh no..

OpenStudy (priyar):

but still.. its not correct right?

imqwerty (imqwerty):

its coming alright tho? \((x-1)^3+2(x-1)^2 = x^3-1-3x^2+3x +2x^2+2-4x\) =\(x^3-x^2-x+1\)

OpenStudy (priyar):

yeah its correct! sry..

imqwerty (imqwerty):

its okay =] after this conversion we can apply binomial and cases reduce in this method

OpenStudy (priyar):

yes..

OpenStudy (priyar):

Thanks!

imqwerty (imqwerty):

np :)